OpenStudy (anonymous):

Given two points on the line, input the equation of the line in standard form Ax + By = C. Reduce all fractional answers to lowest terms. (7, -3), (4, -8)

OpenStudy (anonymous):

First you have to find the slope using the formula : slope(m) = (y2 - y1) / (x2 - x1)... set 1 (7,-3) set 2 (4,-8) m = (-8 - (-3) / (4 - 7) m = (-8 + 3) / (-3) m = - 5/-3 = 5/3 slope is 5/3 now we will use this formula : y - y1 = m(x - x1)....using either of the points, I will use (7,-3) and slope 5/3 y - (-3) = 5/3(x - 7) y + 3 = 5/3x - 35/3 y = 5/3x - 35/3 - 9/3 y = 5/3x - 44/3 Now we need to turn this into standard form Ax + By = C y = 5/3x - 44/3 (subtract 5/3x from both sides) - 5/3x + y = -44/3 (multiply both sides by -3) 5x - 3y = 44 <-- your answer